UROP Proposal

Observation of Koronis Family Asteroids for indications of Collisional Evolution

May 9, 1993

I will be working with Doctor Binzel on observation and data analysis for the Koronis asteroid project. I will be doing observations at Wallace observatory with Steve Slivan, the graduate student on the project. I will be learning observation techniques at the beginning of the summer and will soon be able to observation on my own. The primary observational goal for my summer work will be to continue to collect light curves at multiple oppositions for the Koronis asteroids. Additionally, I will be doing data analysis in an attempt to resolve polar orientation of these asteroids based on the light curves collected from observation. I may also be doing other tasks related to the Koronis project to be determined during the course of the summer by Dr. Binzel and Mr. Slivan.